Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

A group of pictures of butterflies contains 10 pictures

Expert replies
by vishalpathak » Wed Nov 20, 2013 6:12 am
A group of pictures of butterflies contains 10 pictures. Jim had bought 3 of them. If 2 pictures are to be picked out from the 10 pictures, what is the probability that both pictures are not those that were already bought by Jim?

14/15
3/5
6/13
7/15
7/10

[spoiler]OA D
I did: Total number of ways to pick 2: 10C2 = 45
Total number of ways to pick the 2 of Jim's 3: 3C2 = 3
Therefore probability of picking the others = 1 - 3/45 = 42/45 = 14/15. Where did I go wrong?[/spoiler]
Join the discussion
Source: — Problem Solving |

by ganeshrkamath » Wed Nov 20, 2013 6:23 am
vishalpathak wrote:A group of pictures of butterflies contains 10 pictures. Jim had bought 3 of them. If 2 pictures are to be picked out from the 10 pictures, what is the probability that both pictures are not those that were already bought by Jim?

14/15
3/5
6/13
7/15
7/10

[spoiler]OA D
I did: Total number of ways to pick 2: 10C2 = 45
Total number of ways to pick the 2 of Jim's 3: 3C2 = 3
Therefore probability of picking the others = 1 - 3/45 = 42/45 = 14/15. Where did I go wrong?[/spoiler]
Favorable cases:
Select 2 out of 7 not selected by Jim = 7C2

Sample space:
Select 2 out of any of the 10 pictures = 10C2

Probability = 7C2/10C2
= (7*6)/(10*9)
= 7/15

Choose D

Cheers
Every job is a self-portrait of the person who did it. Autograph your work with excellence.

Kelley School of Business (Class of 2016)
GMAT Score: 750 V40 Q51 AWA 5 IR 8
https://www.beatthegmat.com/first-attemp ... tml#688494
Join the discussion

by ganeshrkamath » Wed Nov 20, 2013 6:26 am
vishalpathak wrote:A group of pictures of butterflies contains 10 pictures. Jim had bought 3 of them. If 2 pictures are to be picked out from the 10 pictures, what is the probability that both pictures are not those that were already bought by Jim?

14/15
3/5
6/13
7/15
7/10

[spoiler]OA D
I did: Total number of ways to pick 2: 10C2 = 45
Total number of ways to pick the 2 of Jim's 3: 3C2 = 3
Therefore probability of picking the others = 1 - 3/45 = 42/45 = 14/15. Where did I go wrong?[/spoiler]
I think you are not including some cases:
Total number of ways of picking 1 out of Jim's and one not: 3C1 * 7C1 = 3*7 = 21
Therefore probability of picking the others = 1 - 3/45 - 21/45 = 21/45 = 7/15

Hope this helps.
Cheers
Every job is a self-portrait of the person who did it. Autograph your work with excellence.

Kelley School of Business (Class of 2016)
GMAT Score: 750 V40 Q51 AWA 5 IR 8
https://www.beatthegmat.com/first-attemp ... tml#688494
Join the discussion

by vishalpathak » Wed Nov 20, 2013 6:42 am
ganeshrkamath wrote:
vishalpathak wrote:A group of pictures of butterflies contains 10 pictures. Jim had bought 3 of them. If 2 pictures are to be picked out from the 10 pictures, what is the probability that both pictures are not those that were already bought by Jim?

14/15
3/5
6/13
7/15
7/10

[spoiler]OA D
I did: Total number of ways to pick 2: 10C2 = 45
Total number of ways to pick the 2 of Jim's 3: 3C2 = 3
Therefore probability of picking the others = 1 - 3/45 = 42/45 = 14/15. Where did I go wrong?[/spoiler]
Favorable cases:
Select 2 out of 7 not selected by Jim = 7C2

Sample space:
Select 2 out of any of the 10 pictures = 10C2

Probability = 7C2/10C2
= (7*6)/(10*9)
= 7/15

Choose D

Cheers
got it .. thanks
Join the discussion

by Brent@GMATPrepNow » Wed Nov 20, 2013 6:59 am
vishalpathak wrote:A group of pictures of butterflies contains 10 pictures. Jim had bought 3 of them. If 2 pictures are to be picked out from the 10 pictures, what is the probability that both pictures are not those that were already bought by Jim?

14/15
3/5
6/13
7/15
7/10
The wording of this question is somewhat ambiguous.

If we say that "both pictures are not those that were already bought by Jim," does this mean that NEITHER picture can be pre-purchased, or does it mean that it's okay for 1 picture to be pre-purchased (but not both)?

For example, if I forbid a child to eat BOTH of the candies in the bowl, does this mean that the child can eat one of the candies? Or does it mean that the child can eat NEITHER?

vishalpathak's answer of [spoiler]14/15[/spoiler] is correct if we say that ONE (but not both) of the pictures can be pre-purchased.

ganeshrkamath's answer of [spoiler]7/15[/spoiler] is correct if we say that NEITHER of the pictures can be pre-purchased.

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by [email protected] » Wed Nov 20, 2013 11:40 am
Hi vishalpathak,

ganeshrkamath has one way to do the "math" behind this question. Here's another.

Total Pictures = 10
Jim's = 3
Not Jim's = 7

Probability of picking 2 pictures that are NOT Jim's:

Probability of 1st: 7/10
Probability of 2nd: 6/9

(7/10)(6/9) = 42/90 = 7/15

Final Answer: D

GMAT assassins aren't born, they're made,
Rich
Contact Rich at [email protected]
Image
Join the discussion